### v2

@@ 3.1 Governing equations and boundary conditions
Emulsions are structured two-phase fluids composed of droplets of density $\rho+\Delta\rho$ and viscosity $\lambda\mu$ suspended in a continuous-phase fluid of density $\rho$ and viscosity $\mu$. If both the dispersed and the continuous phase are Newtonian, incompressible fluids, and interface is also Newtonian and slip or dissipation free, the only additional material parameter included is the interfacial tension that depends on the two liquids chosen. Assuming that variations of an emulsion macroscopic flow occur over a characteristic length scale $L$, the linear momentum and mass conservation equations in the continuum limit, and in the absence of body-force torques, are
@@
where $\mathbf{u}$ is the velocity field averaged over a continuum volume of fluid, $\mathbf{\Sigma}$ is volume-averaged stress tensor in the emulsion, and $Re$ is the macroscopic Reynolds number. Although, both suspended- and continuous fluids are typically Newtonian, macroscopic rheological behavior is non-Newtonian (e.g., shear thinning and normal stress differences) due to the interplay of droplet-level deformation and relaxation, interfacial dynamics, and interdrop interactions leading to an anisotropic emulsion microstruture in response to imposed bulk stresses [87].
In most applications where emulsions play a key role, droplet size is within the micron scale or smaller such that the local Reynolds number defined in terms of the local shear rate and particle size is $Re_{local}=Re(a/L)^{2}$ provided that $a/L\ll 1$, where $a$ is the average, undisturbed droplet size. Hence, the dynamics at the droplet level are governed by the low-Reynolds-number flow equations,
@@
where the primes denote quantities associated with the drop phase, $\mathbf{g}$ is the gravitational acceleration, and $p$ is the mechanical pressure. Equations (2)-(3) are valid everywhere expect at the droplet interface denoted by $S$. Often models assume that the suspending liquid is density matched with the droplet or dispersed phase. Boundary conditions encompass, typically, an imposed flow field
@@
where [.$[.]_{S}$ denotes a jump of the bracketed quantity across the interface, $\nabla_{S}=(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\nabla$ is the surface gradient operator,
@@
is the mean curvature, and $\sigma$ is the interfacial tension coefficient which may vary along the droplet interface in response to gradients in temperature or in-homogeneous distribution of surfactant molecules. In such cases, an equation of state and an evolution equation for surfactant concentration are needed for closure [88, 89]. Typically, a relation between surfactant concentration and surface tension coefficient is given by the non-linear Langmuir equation of state [90],
@@
where $\Gamma$ is the surfactant concentration along the interface, $\sigma_{0}$ is the surface tension of the clean (surfactant-free) interface, $R$ is the ideal gas constant, $T$ is the absolute temperature, and $\Gamma_{\infty}$ is the maximum packing concentration of surfactant molecules in a monolayer. In the absence of flow and after surfactant adsorption occurs for a sufficient time, there is an equilibrium surface tension $\sigma_{eq}$ at which the equilibrium surface pressure $\Pi_{eq}=\sigma_{0}-\sigma_{eq}$ is defined for a given equilibrium surfactant concentration, $\Gamma_{eq}$ [91]. The ratio $\Gamma_{eq}/\Gamma_{\infty}$ known as surface coverage indicates the initial fraction of the interface covered with surfactants. Several adsorption isotherms can be used to model surface tension variations in the presence of surfactants. We direct the interested reader to Table 1 of Ref. [90] for a comprehensive list.
@@
3.2 Relevant physicochemical parameters and dimensionless groups
A characteristic length scale for describing deformation, breakup or coalescence of drops, is the underformed drop size, $a$. The capillary relaxation time defined as $\tau_{\sigma}=\mu a/\sigma_{eq}$, provides a possible characteristic time scale. The characteristic time scale for $\lambda\gg 1$ is defined as $\tau_{\sigma}=\lambda\mu a/\sigma_{eq}$ as the larger of the two viscosities determines the time period for shape relaxation.[9, 30] Assuming a neutrally-buoyant drop ($\Delta\rho=0$) in an imposed linear flow field where $\mathbf{u}^{\infty}\sim\mathbf{x}\cdot\nabla\mathbf{u}$, the characteristic time scale for the flow is $\tau_{f}=\dot{\gamma}^{-1}$, where $\dot{\gamma}$ is the magnitude of the local velocity gradient. The $\tau_{\sigma}$ or viscocapillary time captures the time required to traverse a distance comparable to drop size, with an intrinsic capillary velocity, $\sigma_{eq}/\mu$ set by the ratio of two physicochemical parameters or material properties: interfacial tension and the viscosity.[34] The two material quantities can be used for estimating characteristic scale for pressures or stresses, as follows. The ratio $\sigma_{eq}/a$, provides an estimate for capillary stress. A typical timescale for droplet deformation in shear is $\tau_{d}\sim\tau_{f}=\dot{\gamma}^{-1}$. Setting the undeformed drop size, $a$, as the characteristic length scale, a natural choice for the characteristic velocity is $\dot{\gamma}a$ and hence, from Eqs. (2)-(3) the pressures inside and outside of the droplet scale as $\mu\dot{\gamma}$ and $\lambda\mu\dot{\gamma}$, respectively. The choices of characteristic time, length and stress/pressure scales determine the form of dimensionless equations and boundary conditions obtained after a nondimensionalization of Eqs. (2)-(10).
@@ 3.2 Relevant physicochemical parameters and dimensionless groups
Alternatively, $Ca$ equals the ratio of imposed flow velocity, $\dot{\gamma}a$ to intrinsic capillary velocity, $\sigma_{eq}/\mu$. Capillary number can be equivalently written as ratio of capillary relaxation time to deformation time. Since $Ca$ is also a product of relaxation time, $\tau_{\sigma}$ and deformation rate ($\dot{\gamma}$ for shear), it captures the flow strength in a fashion reminiscent of Weissenberg number $Wi=\dot{\gamma}\tau_{1}$ used in polymer rheology, with $\tau_{1}$ representing the longest relaxation time. Thus, $Ca$ captures the relative magnitude of stress, velocity, and flow strength for calibrating the influence of applied flow conditions on drop deformation and dynamics. Again, for $\lambda\gg 1$, the $Ca$ values should be computed by considering $\tau_{\sigma}=\lambda\mu a/\sigma_{eq}$ as the shape relaxation time [9, 30].
Two more dimensionless groups written as the ratio of stresses or pressures. The Bond number, $Bo$ captures the ratio of hydrostatic to capillary pressures, relevant to determining bouyancy-driven motion and gravity-induced drop deformation. surface tension gradients. The the Marangoni number, $Ma$, is a ratio between distorting viscous stresses and restoring Marangoni stresses that arise due to surface tension variation.
$Bo=\frac{\Delta\rho ga}{\sigma_{eq}/a}\,,\quad Ma^{-1}=\frac{\mu\dot{\gamma}}{\Delta\sigma/a}\,,$ (12)
If the origin of the Marangoni stress is a non-uniform surfactant contribution, then the characteristic magnitude of surface-tension variations equals the magnitude of surface compression modulus $\Delta\sigma=-\Gamma_{eq}\left(\partial\sigma/\partial\Gamma\right)_{\Gamma=\Gamma_{eq}}$ that arises from perturbations about the equilibrium surface concentration, $\Gamma_{eq}$ The dimensionless ratio of $\Delta\sigma$ to ${\sigma}_{eq}$ represented by $\beta$ is therefore a surface elasticity parameter [69, 72].
@@
Lastly, $Pe_{S}$ is the surface Péclet number denoting the relative balance between surfactant convection and diffusion along the interface. The process of emulsification by mechanical methods can sometimes require evaluation of inertial effects using the characteristic inertial pressure estimated as ${\rho}U^{2}$. For example, Reynolds number, $Re$ and Weber number, $We$ are defined as the ratio of inertial pressure to viscous and capillary stress, respectively [34].
Dissipative effects stemming from the surface viscosities may affect the dynamics of droplets in flows. Here, a balance between bulk viscous stresses and dissipative interfacial stresses are embedded in two dimensionless Boussinesq numbers,
@@
for shear surface viscosity, $\mu_{s}$, and dilational viscosities, $\mu_{d}$, respectively. In such cases, the right-hand side of the traction jump boundary condition in Eq. (6) is augmented by an additive interfacial-viscous traction of the form, $\nabla_{S}\cdot\tau_{S}$, obeying the deviatoric part of the Boussinesq-Scriven constitutive law for Newtonian interfaces [85, 92, 93, 17],
@@
Accounting for surface viscosity alters the interfacial force balance (6) and affects the interfacial transport of surface-active entities on complex interfaces. Gradients in surface tension $\nabla_{S}\sigma$ generate Marangoni stresses that act to immobilize surfactant transport, while surface viscous stresses oppose surface velocity gradients, where shear viscosity has a similar role as Marangoni stresses and surface dilatational viscosity reduces local surfactant concentration by diluation effects [94, 95, 55, 96]. Recent works suggest that surface viscosity depends exponentially on surface pressure [97, 98, 99, 100, 90]
@@
where $\Pi=\sigma_{0}-\sigma$ is surface pressure, $i=s,d$ stands for shear and dilatational viscosities, $\mu_{i,eq}$ and $\Pi_{eq}$ are the equilibrium surface viscosity and surface pressure, respectively, and $\Pi_{c}$ is a characteristic scale of surface pressure variations. Positive values of $\Pi_{c}$ indicate $\Pi$-thickening surfactants, while negative values are used for $\Pi$-thinning surfactants. The relation between surfactant transport and surface viscous stresses is given by combining Eq. 8 and 16 yielding surfactant-concentration-dependent Boussinesq numbers,
@@
Emulsions of droplets with slip-boundaries have been used as model system to probe the rheology of emulsions of immiscible polymer blends, where the slip coefficient is defined by the ratio of the interfacial thickness and some isotropic interfacial viscosity [54, 101, 102]. Non-dimensionalizing Eq. (5) yields a dimensionless slip coefficient $\bar{\alpha}=\alpha/(\mu a)$.
@@ 3.3 Emulsion macroscopic stress
The continuum, macroscopic volume-averaged stress in Eq. (2) for a particulate system where both dispersed and suspending fluids are Newtonian is
@@
where $\phi$ is the drop-phase volume fraction, $\langle$.$\rangle$ . ⟩ denote the volume-average of the quantity in brackets, $\mathbf{\Sigma}^{0}=-\langle p\rangle\mathbf{I}+2\mu\langle\mathbf{E}\rangle$ is the Newtonian stress contribution from the continuous phase, and
@@
is the particle extra stress in a suspension; the sum accounts for the stress contribution of each one of the $N$ particles in suspension given by
@@
known as the Landau-Batchelor tensor [103], which depends on the surface traction and the velocity distribution over the particle surface, where local low-Reynolds number conditions hold and no external torques are applied. In the limit of a sharp, fluid interface, $\mathbf{\Sigma}\cdot\mathbf{n}\to\Delta\mathbf{f}$ captures the stress jump across the interface defined in Eq. (6). For example, considering clean, neutrally buoyant droplets, $\Delta\mathbf{f}=2H\sigma\mathbf{n}$.
The emulsion shear rheology is defined by a shear viscosity $\Sigma_{12}$ and first- and second-normal stress differences that arise from contributions of the dispersed phase only, [104]
@@
An effective viscosity is derived from the dimensionless form of Eq. (19),
@@
$\bar{\mathbf{\Sigma}}=\frac{\mathbf{\Sigma}}{\mu\dot{\gamma}}\,,\quad\bar{\mathbf{\Sigma}}^{0}=\frac{\mathbf{\Sigma}^{0}}{\mu\dot{\gamma}}\,,\quad\bar{\mathbf{\Sigma}}^{p}=\frac{\mathbf{\Sigma}^{p}}{\sigma_{eq}/a}\,,$ (24)
and $\eta_{r}\equiv(\eta/\mu)=\bar{\Sigma}_{12}$, where $\eta$ is the emulsion continuum viscosity. Even though both drop and suspending fluids are assumed Newtonian, results from small deformation theories and numerical simulations show that emulsions display a significant shear-thinning behavior and finite normal stress differences ($N_{1}>0$ and $N_{2}<0$) indicative of a characteristic non-Newtonian behavior arising from a balance between the material structure relaxation time $\sim\mu a/\sigma_{eq}$ and imposed flow rates $\dot{\gamma}^{-1}$ reminiscent of the Weissenberg number in elastic soft materials. Emulsion drop polydispersity can be included in the derivation of Eq. (18) if the distribution of particle sizes is known. Equations (18)-(23) hold for the analysis of dilute to concentrated suspensions. At higher concentrations, near the maximum volume fraction of drops, more elaborate constitutive equations are needed to adequately probe the material rheology. The state of stress and flow behavior of jammed dense emulsions are discussed in Section 6.

### v3

@@ 3.1 Governing equations and boundary conditions
Emulsions are structured two-phase fluids composed of droplets of density $\rho+\Delta\rho$ and viscosity $\lambda\mu$ suspended in a continuous-phase fluid of density $\rho$ and viscosity $\mu$. If both the dispersed and the continuous phases are Newtonian, incompressible fluids, and interface is also Newtonian and slip or dissipation free, the only additional material parameter needed is the interfacial tension that depends on the two liquids chosen. In the continuum limit, and in the absence of body-force torques, the linear momentum and mass conservation equations are
@@
where $\mathbf{u}$ is the velocity field averaged over a continuum volume of fluid, $\mathbf{\Sigma}$ is volume-averaged stress tensor in the emulsion, and $Re={u}L/\nu$ is the macroscopic Reynolds number defined as ratio of convective to diffusive velocity (or ratio of inertial to viscous stress) [38]. Here we are assuming that variations of an emulsion macroscopic flow occur over a characteristic length scale $L$, such that $a/L\ll 1$, where $a$ is the average, undisturbed droplet size. Even for emulsions with Newtonian flow behavior for both phases, the emulsion exhibit non-Newtonian rheological behavior (e.g., shear thinning and normal stress differences) due to the interplay of droplet-level deformation and relaxation, interfacial dynamics, and interdrop interactions leading to an anisotropic emulsion microstruture in response to imposed bulk stresses [97].
In most applications where emulsions play a key role, droplet size is within the few nanometer to few micron scale such that the local Reynolds number defined in terms of the local shear rate and particle size is $Re_{local}=Re(a/L)^{2}$ provided that $a/L\ll 1$. Hence, the microhydrodynamics at the droplet level are governed by the low-Reynolds-number flow equations,
@@
where the primes denote quantities associated with the drop phase, $\mathbf{g}$ is the gravitational acceleration, and $p$ is the mechanical pressure. Equations (2)-(3) are valid everywhere expect at the droplet interface denoted by $S$. Often models assume that the suspending liquid is density matched with the droplet or dispersed phase.
Boundary conditions encompass, typically, an imposed flow field
@@
where [.$[.]_{S}$ denotes a jump of the bracketed quantity across the interface, $\nabla_{S}=(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\nabla$ is the surface gradient operator, and $\sigma$ is the interfacial tension coefficient which may vary along the droplet interface in response to gradients in temperature or in-homogeneous distribution of surfactant molecules. The mean curvature $H$ is computed using
@@
In such cases, an equation of state and an evolution equation for surfactant concentration, $\Gamma$, are needed for closure [98, 99]. Several adsorption isotherms that outline how surface tension varies with change in interfacial concentration of surfactants can be used. We direct the interested reader to Table 1 of Ref. [100] for a comprehensive list. Here, we illustrate methodology using the non-linear Langmuir equation of state,
@@
where $\sigma_{0}$ is the surface tension of the clean (surfactant-free) interface, $R$ is the ideal gas constant, $T$ is the absolute temperature, and $\Gamma_{\infty}$ is the maximum packing concentration of surfactant molecules in a monolayer. In the absence of flow and after surfactant adsorption occurs for a sufficient time, there is an equilibrium surface tension $\sigma_{eq}$ at which the equilibrium surface pressure $\Pi_{eq}=\sigma_{0}-\sigma_{eq}$ is defined for a given equilibrium surfactant concentration, $\Gamma_{eq}$ [101]. The ratio $\Gamma_{eq}/\Gamma_{\infty}$ known as surface coverage indicates the initial fraction of the interface covered with surfactants.
@@
3.2 Relevant physicochemical parameters, scales, and dimensionless groups
A characteristic length scale for describing deformation, breakup or coalescence of drops, is the undeformed drop size, $a$. A possible characteristic time scale can be defined in terms of the capillary relaxation time or shape relaxation time, written as:
$\tau_{\sigma}=\mu a/\sigma_{eq},\;or\;\tau_{\sigma}=\lambda\mu a/\sigma_{eq},$ (11)
as the larger of the two viscosities determines the time period for shape relaxation, the characteristic time scale for $\lambda\gg 1$ is defined as $\tau_{\sigma}=\lambda\mu a/\sigma_{eq}$ [15, 32]. Otherwise $\tau_{\sigma}=\mu a/\sigma_{eq}$ is typically used, and $\tau_{\sigma}$, also called viscocapillary time, captures the time required to traverse a distance comparable to drop size, with an intrinsic capillary velocity, $\sigma_{eq}/\mu$ set by the ratio of two physicochemical parameters or material properties: interfacial tension and viscosity [38]. The two material parameters can be used to estimate characteristic scale for pressures or stresses, as follows. The ratio $\sigma_{eq}/a$, provides an estimate for capillary stress, whereas $\mu\dot{\gamma}$ estimates the viscous stress.
Assuming a neutrally-buoyant drop ($\Delta\rho=0$) in an imposed linear flow field where $\mathbf{u}^{\infty}\sim\mathbf{x}\cdot\nabla\mathbf{u}$, the characteristic time scale for the flow is $\tau_{f}=\dot{\gamma}^{-1}$, where $\dot{\gamma}$ is the magnitude of the local velocity gradient. A typical timescale for droplet deformation in shear is $\tau_{d}\sim\tau_{f}=\dot{\gamma}^{-1}$. Setting the undeformed drop size, $a$, as the characteristic length scale, a natural choice for the characteristic velocity is $\dot{\gamma}a$ and hence, from Eqs. (2)-(3) the pressures inside and outside of the droplet scale as $\mu\dot{\gamma}$ and $\lambda\mu\dot{\gamma}$, respectively. The choices of characteristic time, length and stress/pressure scales determine the form of dimensionless equations and boundary conditions obtained after a nondimensionalization of Eqs. (2)-(10).
@@ 3.2 Relevant physicochemical parameters, scales, and dimensionless groups
Alternatively, $Ca$ equals the ratio of imposed flow velocity, $\dot{\gamma}a$ to intrinsic capillary velocity, $\sigma_{eq}/\mu$. The capillary number can be written equivalently as the ratio of capillary relaxation time to deformation time. Since $Ca$ is also a product of relaxation time, $\tau_{\sigma}$ and deformation rate ($\dot{\gamma}$ for shear), it captures the flow strength in a fashion reminiscent of Weissenberg number $Wi=\dot{\gamma}\tau_{1}$ used in polymer rheology, with $\tau_{1}$ representing the longest relaxation time. Thus, $Ca$ captures the relative magnitude of stress, velocity, and flow strength for calibrating the influence of applied flow conditions on drop deformation and dynamics. Again, for $\lambda\gg 1$, the $Ca$ values should be computed by considering $\tau_{\sigma}=\lambda\mu a/\sigma_{eq}$ as the shape relaxation time [15, 32].
Two additional dimensionless groups are written as the ratio of stresses or pressures. The Bond number, $Bo$ captures the ratio of hydrostatic to capillary pressures, relevant to determining buoyancy-driven motion and the influence of gravity on the shape and deformation of drops. The Marangoni number, $Ma$, is a ratio between restoring Marangoni stresses $\Delta\sigma/a$ that arise due to surface tension variation, $\Delta\sigma$ and distorting viscous stresses,
$Bo=\frac{\Delta\rho ga}{\sigma_{eq}/a}\,,\quad Ma^{-1}=\frac{\mu\dot{\gamma}}{\Delta\sigma/a}\,.$ (13)
If the origin of the Marangoni stress is a non-uniform surfactant contribution, then the characteristic magnitude of surface-tension variation equals the magnitude of surface compression modulus $\Delta\sigma=-\Gamma_{eq}\left(\partial\sigma/\partial\Gamma\right)_{\Gamma=\Gamma_{eq}}$ that arises from perturbations about the equilibrium surface concentration, $\Gamma_{eq}$. The dimensionless ratio of $\Delta\sigma$ to ${\sigma}_{eq}$ represented by $\beta$ is a surface elasticity parameter [74, 77],
@@
where, $Pe_{S}$ is the surface Péclet number denoting the relative balance between surfactant convection and diffusion along the interface. Modeling emulsification by mechanical methods [33, 34, 103, 104] can sometimes require the evaluation of inertial effects using the characteristic inertial pressure estimated as ${\rho}U^{2}$. For example, Reynolds number, $Re=\rho U^{2}/(\mu U/a)$ and Weber number, $We=\rho U^{2}/(\sigma/a)$ are defined as the ratio of inertial pressure to viscous and capillary stress, respectively [38].
Dissipative effects due shear and dilatational surface viscosity may affect the dynamics of droplets in flows [20]. Two dimensionless Boussinesq numbers that capture the interplay between bulk viscous stresses and dissipative interfacial stresses, are defined as
@@
for shear surface viscosity, $\mu_{s}$, and dilatational viscosity, $\mu_{d}$, respectively. In such cases, the right-hand side of the traction jump boundary condition in Eq. (6) is augmented by an additive interfacial-viscous traction of form, $\nabla_{S}\cdot\tau_{S}$, obeying the deviatoric part of the Boussinesq-Scriven constitutive law for Newtonian interfaces [95, 105, 106, 19],
@@
Incorporating surface viscosity can alter the interfacial force balance in Eq. (6) and interfacial transport of surface-active entities at complex interfaces. Gradients in surface tension ($\nabla_{S}\sigma$) generate Marangoni stresses. Interfacial shear viscosity characterizes the resistance to interfacial shear flow, and surface dilatational viscosity captures the resistance to dilatational effects that can influence coalescence [107, 108, 60, 109]. Interfacial concentration and interaction between adsorbed molecules (and macromolecules) influence interfacial tension, $\sigma$ and surface pressure, $\Pi=\sigma_{0}-\sigma$ that depends on the surface tension reduction compared to value at a clean interface, $\sigma_{0}$.
Several studies suggest that surface viscosity depends exponentially on surface pressure [110, 111, 112, 113, 100]
@@
where $i=s,d$ identify shear and dilatational viscosities, $\mu_{i,eq}$ and $\Pi_{eq}$ are the equilibrium surface viscosity and surface pressure, respectively, and $\Pi_{c}$ is a characteristic scale of surface pressure variations. Positive values of $\Pi_{c}$ indicate $\Pi$-thickening surfactants, while negative values are used for $\Pi$-thinning surfactants. The relation between surfactant transport and surface viscous stresses is given by combining Eqs. (8) and (17) yielding the surfactant-concentration-dependent Boussinesq numbers, defined as follows
@@
Emulsions of droplets with slip-boundaries have been used to model the rheology of emulsions of immiscible polymer blends, where the slip coefficient is defined by the ratio of the interfacial thickness and some isotropic interfacial viscosity [59, 114, 115]. Non-dimensionalizing Eq. (5) yields a dimensionless slip coefficient $\bar{\alpha}=\alpha/(\mu a)$.
Emulsions with one or both phases as non-Newtonian require additional parameters and considerations, which depend on the choice of constitutive model made for capturing one or more features typical of non-Newtonian behavior: rate-dependent shear and extensional viscosity, first and second normal stress difference, and relaxation time. Even in the simplest case of the second-order fluid model for both phases, two normal stress differences ${N_{1i}}$ and ${N_{2i}}$ each arise for dispersed ($i=d$) and suspending ($i=s$) phase, creating at least four additional dimensionless parameters:
${N_{1i}}a/\sigma_{eq},\>N_{2s}/N_{1s},\>N_{2d}/N_{1d},\>De_{i}=\tau_{i}/\tau_{\sigma},$ (19)
where $\tau_{i}={N_{1i}/\Sigma_{12i}}$ can be used for defining the relaxation time for suspending or disperse phase, in which case, if we define $\bar{N}_{1i}={N_{1i}}a/\sigma_{eq}$ then $De_{i}=\bar{N}_{1i}/{Ca}^{2}$, [116, 9, 45]. The relaxation time for viscoelastic fluid phase can be alternatively determined using the linear viscoelastic response measured in oscillatory shear, stress relaxation, dynamic light scattering or capillarity-based extensional rheology, and each response captures aspects of non-Newtonian response that need not correlate directly with the first normal stress difference.
@@ 3.3 Emulsion macroscopic stress
The continuum, macroscopic volume-averaged stress in Eq. (2) for an emulsion where both dispersed and suspending fluids are Newtonian is
@@
where $\phi$ is the drop-phase volume fraction, and $\langle$.$\rangle$ . ⟩ denote the volume-average of the quantity in brackets. Here $\mathbf{\Sigma}^{0}=-\langle p\rangle\mathbf{I}+2\mu\langle\mathbf{E}\rangle$ is the Newtonian stress contribution from the continuous phase. In analogy with a particulate system [117], the extra stress in an emulsion due to the dispersed droplets can be determined using the following expression:
@@
where the sum accounts for the stress contribution of each one of the $N$ drops in emulsion (or particles in suspension) given by the Landau-Batchelor tensor [118] defined as
@@
The Landau-Batchelor tensor depends on the surface traction and the velocity distribution over the particle surface, where local low-Reynolds number conditions hold and no external torques are applied. In the limit of a sharp fluid interface, $\mathbf{\Sigma}\cdot\mathbf{n}\to\Delta\mathbf{f}$ captures the stress jump across the interface defined in Eq. (6). For example, considering clean, neutrally buoyant droplets, $\Delta\mathbf{f}=2H\sigma\mathbf{n}$.
Thus, the connection between microscopic behavior and macroscopic rheology is embedded in the definition of the macroscopic particle-stress contribution $\mathbf{\Sigma}^{p}$ given by equation Eq. (21). The stress jump across the interface, $\Delta\mathbf{f}$, captures the microscale physics while $\mathbf{\Sigma}^{p}$ accounts for the contribution of the dispersed phase. The emulsion shear rheology is defined by a shear viscosity $\Sigma_{12}$ and first- and second-normal stress differences that arise from contributions of the dispersed phase only, [119]
@@
Relative viscosity, $\eta_{r}\equiv(\eta/\mu)$ that equals emulsion viscosity $\eta$ scaled by the suspending fluid viscosity $\mu$ can be defined in terms of $\bar{\Sigma}_{12}$the dimensionless form of extra stress due to added particles or drops as given by Eq. (21). Thus, the relative viscosity has the following form
@@
$\bar{\mathbf{\Sigma}}=\frac{\mathbf{\Sigma}}{\mu\dot{\gamma}}\,,\quad\bar{\mathbf{\Sigma}}^{0}=\frac{\mathbf{\Sigma}^{0}}{\mu\dot{\gamma}}\,,\quad\bar{\mathbf{\Sigma}}^{p}=\frac{\mathbf{\Sigma}^{p}}{\sigma_{eq}/a}\,.$ (26)
Even though both dispersed and suspending fluids are assumed Newtonian, experimental, theoretical and numerical simulations show that emulsions can exhibit non-Newtonian response, including shear-thinning and finite normal stress differences ($N_{1}>0$ and $N_{2}<0$). The surface-tension-driven recovery or relaxation of the perturbed drop shape to the minimum surface area for a fixed volume underlies the origin of viscoelasticity. The ratio of this shape relaxation time $\sim\mu a/\sigma_{eq}$ to imposed flow rates $\dot{\gamma}^{-1}$ is a dimensionless group, defined as $Ca$ in Eq. (12) that evokes the Weissenberg number, $Wi$ in elastic soft materials, and likewise, a nonlinear response is observed for $Wi$ > 1. Equations (20)-(25) hold for the analysis of dilute to concentrated suspensions. Polydispersity in drop sizes of emulsions can be included in the derivation of Eq. (20) if the distribution of drop sizes is known. At higher concentrations, near the maximum volume fraction of drops, more elaborate constitutive equations are needed to adequately capture the rheological response of the emulsions. The stress and flow behavior of the jammed dense emulsions are discussed in Section 6.